# SERA-H: Beyond Native Sentinel Spatial Limits for High-Resolution Canopy Height Mapping

### v2

@@ 4 Discussion
These results position SERA-H as a competitive alternative to approaches based on commercial higher-resolution imagery. Although Fogel obtained slightly sharper boundaries through the use of higher-resolution satellite images, SERA-H offers statistically comparable accuracy using only Sentinel-1 and Sentinel-2 data, which is freely available worldwide. Furthermore, while SPOT images are generally available on an annual basis, the Sentinel-1 and Sentinel-2 satellites offer a significantly higher revisit frequency. This temporal density allows estimates to be updated several times during the year, overcoming the constraint of static annual mapping and paving the way for possible intra-annual forest monitoring using freely available images.
@@
5 Conclusion and Limitations
Model Input Reference MAE RMSE nMAE Tree
images images (m) (m) (%) cover IoU (%)
@@ 5 Conclusion and Limitations
In this paper, we introduced SERA-H, a new end-to-end model combining super-resolution and spatio-temporal learning to generate very high-resolution (2.5 m) canopy height maps from coarser (10 m) Sentinel-1 and Sentinel-2 imagery. Our methodology was validated through a comprehensive ablation study, which confirmed the critical role of the super-resolution module (EDSR) and the temporal attention mechanism (UTAE) in reconstructing fine canopy details that standard interpolation methods fail to capture. Furthermore, we benchmarked SERA-H against five recent state-of-the-art methods, ranging from Sentinel-based models trained on GEDI (Schwartz et al., 2025; Pauls et al., 2025) to approaches leveraging commercial very high-resolution imagery supervised by ALS data (Fogel et al., 2025; Tolan et al., 2024; Liu et al., 2023). The results demonstrate that, by leveraging the temporal density of open-access data, SERA-H not only outperforms other Sentinel-based models but also achieves accuracy comparable to, and in some cases superior to, models relying on costly commercial imagery. This effectively decouples prediction accuracy from input resolution, positioning SERA-H as a potential precise and cost-effective tool for forest management and biodiversity monitoring.
@@
Despite these significant advances, the method is subject to intrinsic physical and methodological limitations. While SERA-H significantly refines the predicted canopy structure, it remains constrained by the spectral information available in 10 m input pixels; consequently, it cannot fully reconstruct small isolated trees or fine details that are spectrally indistinguishable at this resolution, whereas native very high-resolution imagery might still capture these features. Additionally, the model’s performance is currently linked to the availability of high-density ALS data for supervision. Future work should therefore focus on evaluating the model’s generalization capabilities in data-scarce regions, such as tropical or boreal biomes. To mitigate the dependence on dense LiDAR campaigns, exploring transfer learning or weakly supervised techniques constitutes a promising direction for extending this approach to other areas.
Acknowledgements
@@
We are grateful to Jean-Pierre Renaud from the Office National des Forêts (ONF) for providing the 2020 Vosges ALS dataset. This work was supported by the One Forest Vision Initiative (OFVi) and the AI4Forest project (ANR-22-FAI1-0002). We also acknowledge support from the Danish National Research Foundation through the Center for Remote Sensing and Deep Learning of Global Tree Resources (TreeSense, DNRF192). This work was performed using HPC resources from GENCI-IDRIS (Grants 2025-AD010114718R1 and 2026-AD010114718R2).
Data and Code Availability
@@
The Open-Canopy dataset used in this study is available on Hugging Face at https://huggingface.co/datasets/A14Forest/Open-Canopy.
@@ Appendix A Evaluation at Standardized 10m Resolution
The quantitative results presented in  Table˜3 confirm that Sentinel-2-based methods, such as Pauls and Schwartz, exhibit improved performance when evaluated at their native resolution of 10 m compared to finer scales. By avoiding upsampling artifacts, their Mean Absolute Errors drop to $3.72$ m and $3.36$ m, respectively.
Figure 8: Evolution of the Mean Absolute Error (MAE) of SERA-H as a function of the number of initial image triplets (orange curve, left axis). The grey bars represent the proportion of samples in the test dataset for each bin (right axis). The number of triplets refers to the number of unique acquisitions available before the duplication or truncation process used to reach the fixed input sequence length of 16 required by SERA-H
@@ Appendix B Impact of Temporal Density on Model Performance
However, despite this more favorable evaluation setting for standard Sentinel-2 baselines, SERA-H maintains a significant performance lead across the dataset, achieving an MAE of $3.00$ m. This superiority is particularly evident in the spatial accuracy of the predictions, as reflected by the Tree Cover Intersection over Union (IoU). While Pauls and Schwartz achieve IoU scores of approximately 5555%, suggesting difficulties in precisely delineating canopy boundaries, SERA-H reaches a Tree Cover IoU of 89.3589.35%. This result highlights our model’s ability to correctly localize vegetation, rivalling the performance of Fogel’s high-resolution method (89.9589.95%) and demonstrating that SERA-H successfully reconstructs fine spatial details even from coarser input data.
To assess the robustness of SERA-H regarding data availability, we analyzed the Mean Absolute Error (MAE) as a function of the number of available image triplets ($S2,S1_{asc},S1_{dsc}$) before the duplication or truncation step.

### v3

@@ 4 Discussion
These results position SERA-H as a competitive alternative to approaches based on commercial higher-resolution imagery. Although Fogel obtained slightly sharper boundaries through the use of higher-resolution satellite images, SERA-H offers statistically close accuracy using only Sentinel-1 and Sentinel-2 data, which is freely available worldwide. Furthermore, while SPOT images are generally available on an annual basis, the Sentinel-1 and Sentinel-2 satellites offer a significantly higher revisit frequency. This temporal density allows estimates to be updated several times during the year, overcoming the constraint of static annual mapping and paving the way for possible intra-annual forest monitoring using freely available images.
@@
Model Input Reference MAE RMSE nMAE Tree cover
images CHM (m) (m) (%) IoU (%)
@@ 5 Conclusion and Limitations
In this paper, we introduced SERA-H, a new end-to-end model combining super-resolution and spatio-temporal learning to generate very high-resolution (2.5 m) canopy height maps from coarser (10 m) Sentinel-1 and Sentinel-2 imagery. Our methodology was validated through a comprehensive ablation study, which confirmed the critical role of the super-resolution module (EDSR) and the temporal attention mechanism (UTAE) in reconstructing fine canopy details that standard interpolation methods fail to capture. Furthermore, we benchmarked SERA-H against five recent state-of-the-art methods, ranging from Sentinel-based models trained on GEDI (Schwartz et al., 2025; Pauls et al., 2025) to approaches leveraging commercial very high-resolution imagery supervised by ALS data (Fogel et al., 2025; Tolan et al., 2024; Liu et al., 2023). The results demonstrate that, by leveraging the temporal density of open-access data, SERA-H not only outperforms other Sentinel-based models but also achieves accuracy comparable to, and in some cases superior to, models relying on costly commercial imagery. This effectively decouples prediction accuracy from input resolution, positioning SERA-H as a potentially inexpensive tool for monitoring forest disturbances.
@@
Despite these significant advances, our method is subject to intrinsic physical and methodological limitations. While SERA-H significantly refines the predicted canopy structure, it remains constrained by the spectral information available in 10 m input pixels; consequently, it cannot fully reconstruct small isolated trees or fine details that are spectrally indistinguishable at this resolution, whereas native very high-resolution imagery might still capture these features. Furthermore, SERA-H remains sensitive to signal saturation, a physical constraint common to satellite-based height estimation. Similar to other state-of-the-art models, a performance decline is observed for tall trees, particularly those exceeding 30 m, where the sensitivity of optical and radar signals to canopy height plateaus. Additionally, the model’s performance is currently linked to the availability of high-density ALS data for supervision. Future work should therefore focus on evaluating the model’s generalization capabilities in data-scarce regions, such as tropical or boreal biomes. To mitigate the dependence on dense LiDAR campaigns, exploring transfer learning or weakly supervised techniques constitutes a promising direction for extending this approach to other areas.
Figure 8: Evolution of the Mean Absolute Error (MAE) of SERA-H as a function of the number of initial image triplets (orange curve, left axis). The gray bars in the background represent the proportion of samples in the test dataset for each bin (right axis). The number of triplets refers to the number of unique acquisitions available before the duplication or truncation process used to reach the fixed input sequence length of 16 required by SERA-H
Funding and Acknowledgements
@@
We are grateful to Jean-Pierre Renaud from the Office National des Forêts (ONF) for providing the 2020 Vosges ALS dataset. This work was supported by the One Forest Vision Initiative (OFVi), the AI4Forest project (ANR-22-FAI1-0002) and and the PEPR FORESTT PC Monitor project (ANR-24-PEFO-0003). We also acknowledge support from the Danish National Research Foundation through the Center for Remote Sensing and Deep Learning of Global Tree Resources (TreeSense, DNRF192). This work was performed using HPC resources from GENCI-IDRIS (Grants 2025-AD010114718R1 and 2026-AD010114718R2).
Data Availability
@@
The Open-Canopy dataset used as reference in this study is available on Hugging Face at https://huggingface.co/datasets/A14Forest/Open-Canopy.
@@ Appendix A Evaluation at Standardized 10m Resolution
The quantitative results presented in  Table˜3 confirm that Sentinel-based methods, such as Pauls and Schwartz, exhibit improved performance when evaluated at their native resolution of 10 m compared to finer scales. By avoiding upsampling artifacts, their Mean Absolute Errors drop to $3.72$ m and $3.36$ m, respectively.
However, despite this more favorable evaluation setting for standard Sentinel baselines, SERA-H maintains a significant performance lead across the dataset, achieving an MAE of $3.00$ m. This superiority is particularly evident in the spatial accuracy of the predictions, as reflected by the Tree Cover Intersection over Union (IoU). While Pauls and Schwartz achieve IoU scores of approximately 5555%, suggesting difficulties in precisely delineating canopy boundaries, SERA-H reaches a Tree Cover IoU of 89.3589.35%. This result highlights our model’s ability to correctly localize vegetation, rivalling the performance of Fogel’s high-resolution method (89.9589.95%) and demonstrating that SERA-H successfully reconstructs fine spatial details even from coarser input data.
@@ Appendix B Impact of Temporal Density on Model Performance
To assess the robustness of SERA-H regarding data availability, we analyzed the Mean Absolute Error (MAE) as a function of the number of available image triplets $(S2,S1_{\text{asc}},S1_{\text{dsc}})$ before the duplication or truncation step (see Section˜2.2.1)